OpenStudy (anonymous):

find an equation in slope intercept form for the line through point (-3, 3) and perpendicular to the line containing the two points (2,3) and (1,-2)

OpenStudy (aroub):

-2-3/ 1-2 -5/-1 5 y-y1=m(x-x1) , m= -1/5 , (-3,3) y-3=-1/5(x+3)

OpenStudy (aroub):

sry this was the "point slope form" ^^^^^^

OpenStudy (aroub):

this is the slope intercept form ----> y= -1/5x +6

OpenStudy (anonymous):

y=-(1/5)x+(12/5) i think

OpenStudy (aroub):

y=mx+b , m=-1/5 , (-3,3) 3=-1/5(-3) +b 3=3/5 +b -15/5+9/5=b -6/5=b so the equation is y=-1/5x-6/5 live2 dance thanks for correcting me all the time today!!!!!! :D :D

OpenStudy (anonymous):

no problem :) but in ur last explanation why didnt u subtract the 3/5 from the 3 to get b? (and im not trying 2 be a smarta** im seriously just wondering)